The Wilson's pay their babysitter $5 per hour before 11 pm and $7.50 per hour after 11 pm. one evening they went out for 5 hours and paid the sitter $30.00. what time did they come home?
Answer by stanbon(57214)

The Wilson's pay their babysitter $5 per hour before 11 pm and $7.50 per hour after 11 pm. one evening they went out for 5 hours and paid the sitter $30.00. what time did they come home?
------------------------------------
Let number of hrs. before 11 pm be "x".
Let number of hrs. after 11 pm be "y".
--------------------------
Hours equation: x + y = 5
Value equation: 5x + 7.5y = 30
-------------------------------
x = 5-y
Substitute to get:
5(5-y) + 7.5y = 30
25 -5y +7.5y = 30
2.5y = 5
y = 2 hrs (# of hours after 11:00)
---------------
They arrived home at 1:00 am
=================================
Cheers,
Stan H.